1. You can buy a contract which will pay $ 10,000 per year for 3 years for $26,500. If investments of similar risk return 8% per annum, compounded annually, should you buy the contract?

2. How much do you need to deposit TODAY into an account which earn 8% ANNUAL interest compounded Quarterly so that you will have $ 50,000 in the account seven years from now?

3. What ANNUAL rate of return did an investment pay if a $ 15,250 deposited FIVE YEARS AGO has grown to $ 23,460 TODAY?

4. What is the maximum amount you should be willing to pay TODAY for a contract which will pay $12,500 per year for 5 years beginning one year from today if contracts (investments) of similar risk return 7% per year?
Answer #1

1.
=10000/8%*(1-1/1.08^3)=25770.9698724788
No dont buy at 26500

2.
=50000/(1+8%/4)^(4*7)=28718.727644735

3.
=(23460/15250)^(1/5)-1=8.99626918932275%

4.
=12500/7%*(1-1/1.07^5)=51252.4679493449